Find the number of atoms in a copper rod with a length of 8.15 cm and a radius of 1.15 cm. The density of copper is 8.96 g/cm3.
Answer #1

volume of rod = pi*r^2*h

= 3.14 * (1.15)^2 * (8.15)

= 33.8 cm^3

mass of rod = density of rod * volume of rod

= 8.96 g/cm^3 * 33.8 cm^3

= 302.8 g

Molar mass of Cu = 63.55 g/mol

mass(Cu)= 302.8 g

use:

number of mol of Cu,

n = mass of Cu/molar mass of Cu

=(3.028*10^2 g)/(63.55 g/mol)

= 4.765 mol

use:

number of atoms = number of moles * Avogadro’s number

= 4.765 * 6.022*10^23 atoms

= 2.869*10^24 atoms

Answer: 2.87*10^24 atoms